Grade 12th passElectrostatics

A train 100 m long meets a man going in opposite direction at the speed of 5 km/hr and passes him in 7.2 seconds. At what rate is the train going?

Arun
7 Years agoGrade 12th pass

1 Answer

Vikas TU
7 Years ago
Hiii Arun 
The length of train= 100 meter
Velocity of man = 5 km/hr = 1.39m/s
Let velocity of train = v m/s 
so relative velocity = v+1.39 m/s 
100/7.2 = v+1.39
so v = 12.49 m/s 
v is aproximately equal to 45 km/hr
hope this help.